How do you find the measure of each exterior angle of a polygon? When you have the number of sides, for example, 15?
Found 2 solutions by Alan3354, TutorDelphia:
Answer by Alan3354(69443)

The sum of the Ext angles is 360 for all polygons.
For 15, = 360/15 = 24 degrees

Answer by TutorDelphia(193)

The exterior angles of a polygon add up to 360.
So if you have 15 sides, we can do 15*a=360
So a=24. Each angle is 24 degrees.
http://www.mathopenref.com/polygonexteriorangles.html is a good guide to this concept.